OpenStudy (anonymous):

CALCULUS HELP>> PLEASE! find the critical points for f(x)=x^4-10x^2-6 calculated f'=4x^3-20x factored out 4x to get f'=4x(x^2-5) set x^2-5=0 and got sqrt5 as my answer but its not accepted

OpenStudy (anonymous):

should be 3 critical points \[0,\sqrt{5},-\sqrt{5}\]

OpenStudy (anonymous):

(A) Interval = [-3,-1]. Absolute maximum = Absolute minimum = (B) Interval = [-4,1]. Absolute maximum = Absolute minimum = (C) Interval = [-3,4]. Absolute maximum = Absolute minimum = here is the second part....i found -sqrt5 to be the max(i think) for part A but i guess it is wrong

OpenStudy (calculusfunctions):

Oh OK, now you mention the intervals!

OpenStudy (anonymous):

hehe yes, wanted to make sure i had the basics down

OpenStudy (calculusfunctions):

The critical numbers are: -√5, 0, and √5 (A) [-3, -1] Which of the three critical numbers fall in this interval?

OpenStudy (anonymous):

-sqrt(5) aaand it doesnt accept it

OpenStudy (calculusfunctions):

Right because that is not automatically the answer. find f(-3), f(-√5), and f(-1). The largest of these values will be the absolute maximum, and the smallest of these values will be the absolute minimum. Go ahead!

OpenStudy (anonymous):

yes f(-3) and f(-1) are both -15, f(-sqrt5) is -31 so i guess it would be the min but nonetheless it does not accept......im supposed to plug them into the original equation, correct?

OpenStudy (calculusfunctions):

f(-3) = f(-1) = -15 are the absolute maximum and f(-√5) = -31 is the absolute minimum in the interval [-3, -1]

OpenStudy (anonymous):

oooookay it wanted the values, not the inputs......thanks man !

OpenStudy (calculusfunctions):

Yes! A value is the y value. Get it?

OpenStudy (calculusfunctions):

Now do the rest the same way.

OpenStudy (anonymous):

haha yes, crystal, just didnt read close enough i guess

OpenStudy (calculusfunctions):

Excellent! Need anything else?

OpenStudy (anonymous):

yes sir, have to do the same business on the function 3x-11ln(8x) calculated the derivative to be 3- 11/x but i dont think that is the correct derivative....

OpenStudy (calculusfunctions):

\[\frac{ d(\ln f(x)) }{ dx }=\frac{ f \prime(x) }{ f(x) }\] If\[y =3x -11\ln (8x)\]then\[\frac{ dy }{ dx }=3-\frac{ 88 }{ 8x }\]Thus\[\frac{ dy }{ dx }=3-\frac{ 11 }{ x } \]Therefore @gnarballs you were correct!

OpenStudy (calculusfunctions):

I have to sign out now but i hope that helps!

OpenStudy (anonymous):

setting that =0 would the answer be 3/11? im horrible with fractions

OpenStudy (calculusfunctions):

No, it's x = 11/3

OpenStudy (calculusfunctions):

OK? Is that good because I have to go.

OpenStudy (anonymous):

yes its awesome.....thanks man.....have a good thanksgiving

OpenStudy (calculusfunctions):

I'm from Canada, where we have thanksgiving in October. LOL But thanks and you have yourself a great one also!

OpenStudy (anonymous):

ooohh wuuut aha welp have a good day tomorrow then, thanks again

OpenStudy (calculusfunctions):

Thanks, bye!